# Text-Guided Layer Fusion Mitigates Hallucination in Multimodal LLMs

### v1

@@ 4.2.1 Hallucination mitigation
4.2.2 Fine-grained perception
@@ 4.2.3 General reasoning
4.3 Discussion
4.3.1 Router Layer Selection Dynamics
@@
4.3.2 Impact of Multimodal Guidance
We compare the Text-Only and Multimodal versions of our router. Incorporating a global visual token consistently improves performance on benchmarks requiring precise grounding, such as HallusionBench, POPE, and TextVQA. Visual context helps the router disambiguate whether the query demands global semantics or local fine-grained evidence, enabling more accurate layer reweighting. This validates our hypothesis that multimodal guidance strengthens routing decisions beyond what text alone can provide.
4.3.3 Effect of Load-balancing Loss
@@ 5 Conclusion
In this work, we revisited a core assumption in multimodal large language models: the use of a single deep-layer visual representation as the primary input to the LLM. Our study shows that this design restricts fine-grained grounding and increases the likelihood of hallucination. To address this limitation, we introduced TGIF, a lightweight and training-efficient module that performs text-guided inter-layer fusion over a frozen vision encoder. TGIF treats encoder layers as depth-wise experts and routes them according to the input query, allowing the LLM to access richer, hierarchy-aware visual information without modifying the token budget or updating the vision encoder.
Comprehensive experiments across hallucination benchmarks, OCR tasks, and general VQA confirm that TGIF improves grounding quality and reduces hallucination. The method achieves state-of-the-art results among 7B-scale LLaVA variants and even surpasses larger models on HallusionBench. An analysis of router behaviors shows that TGIF produces semantically meaningful depth-selection patterns: it prioritizes early layers for hallucination detection, mid-level layers for OCR, and more diverse mixtures for open-ended reasoning. This pattern indicates that dynamic fusion provides both adaptivity and interpretability.
Although TGIF substantially improves hallucination robustness and fine-grained perception, several challenges remain. First, our fusion mechanism operates over fixed CLIP features; while dynamic routing improves layer selection, the model remains constrained by the representational biases and resolution limits of the frozen encoder. Future work may explore pairing TGIF with higher-resolution or task-specific encoders, or integrating lightweight vision-side adapters to further enhance fine-grained cues. Second, TGIF routes which layers to use but not what spatial regions within each layer to emphasize. Combining inter-layer fusion with adaptive spatial attention or region-level routing may further strengthen grounding in cluttered or text-heavy scenes.
Overall, TGIF demonstrates that hierarchy-aware visual fusion is a promising and scalable path toward more grounded, trustworthy, and task-aware multimodal LLMs, while also opening opportunities for deeper exploration of dynamic visual representations.
[…]
Supplementary Material
6 Implementation Details
@@
7 Benchmark Descriptions
7.1 Hallucination Evaluation
POPE [Li-hallucination-2023].
@@ 7.2 OCR Evaluation
8 Additional Quantitative Results
8.1 POPE Breakdown
@@
8.2 HallusionBench Leaderboard
Notably, despite utilizing a significantly smaller language backbone (7B parameters), TGIF (49.94%) outperforms several larger open-source baselines, including the 13B-parameter LLaVA-1.5 (46.94%) and the 12.1B-parameter BLIP2-T5 (48.09%) [liu2023improved, Li2023BLIP2BL]. It secures the third rank overall, trailing only the closed-source models GPT-4V and Claude 3 [2023GPT4VisionSC, claude3]. This result shows the efficiency of our text-guided fusion strategy: by dynamically routing to the most relevant visual features, TGIF extracts rich grounding signals from the frozen encoder, effectively mitigating hallucination even against models with nearly double the parameter count.
8.3 OCRBench Performance
Table 5: POPE Results by Subset. Comparison of LLaVA-1.5-7B vs. LLaVA-1.5-7B+TGIF across specific hallucination settings. Best performance per subset is bold.
Subset Method F1 Acc Prec Rec Yes %
@@ 8.3 OCRBench Performance
Table 6: HallusionBench Correctness Leaderboard. We report Question-pair Accuracy (qAcc), Figure Accuracy (fAcc), and All Accuracy (aAcc). Top-3 models under the GPT-4–assisted evaluation are highlighted in bold.
Method Params Eval Mode qAcc $\uparrow$ fAcc $\uparrow$ aAcc $\uparrow$
GPT-4V [2023GPT4VisionSC] (Oct 2023) - Human 31.42 44.22 67.58
@@
Claude 3 [claude3] - GPT-4 21.76 28.61 56.86
Gemini Pro Vision [google_gemini] - GPT-4 7.69 8.67 36.85
BLIP2-T5 [Li2023BLIP2BL] 12.1B GPT-4 15.16 20.52 48.09
Qwen-VL [Bai2023QwenVLAF] 9.6B GPT-4 5.93 6.65 39.15
Open-Flamingo [alayrac2022flamingo] 9B GPT-4 6.37 11.27 38.44
MiniGPT-5 [Zheng2023MiniGPT5IV] 8.2B GPT-4 10.55 9.83 40.30
MiniGPT-4 [zhu2023minigpt] 8.2B GPT-4 8.79 10.12 35.78
InstructBLIP [dai2023instructblip] 8.2B GPT-4 9.45 10.11 45.26
BLIP-2 [Li2023BLIP2BL] 8.2B GPT-4 5.05 12.43 40.48
mPLUG-Owl v2 [ye2023mplugowl2] 8.2B GPT-4 13.85 19.94 47.30
mPLUG-Owl v1 [ye2023mplug] 7.2B GPT-4 9.45 10.40 43.93
LRV-Instruction [liu2023aligning] 7.2B GPT-4 8.79 13.01 42.78
TGIF (Ours) 7B GPT-4 17.36 23.70 49.94
GIT [Wang2022GITAG] 0.8B GPT-4 5.27 6.36 34.37
Random Chance - GPT-4 15.60 18.21 45.96
Table 7: Detailed Results on OCRBench. Breakdown of sub-tasks: Text Recognition (Recog.), Scene Text VQA ($VQA^{S}$), Document VQA ($VQA^{D}$), Key Information Extraction (KIE), and Handwritten Math (HMER). Best results are marked in bold.
@@
Gemini Pro [google_gemini] 215 174 128 134 8 659
GPT-4V [2023GPT4VisionSC] 167 163 146 160 9 645
mPLUG-Owl2 [ye2023mplugowl2] 153 153 41 19 0 366
LLaVAR [Zhang2023LLaVAREV] 186 122 25 13 0 346
LLaVA-1.5-13B [liu2023improved] 176 129 19 7 0 331
LLaVA-1.5-7B [liu2023improved] 160 117 15 5 0 297
TGIF (Ours) 162 121 24 6 0 313
mPLUG-Owl [mplug-owl] 172 104 18 3 0 297
InstructBLIP [instructblip] 168 93 14 1 0 276
BLIP-2 [Li2023BLIP2BL] 154 71 10 0 0 235
MiniGPT-4 v2[minigpt-4] 124 29 4 0 0 157

### v2

@@ 4.3.1  POPE and HallusionBench
4.3.2  Trade-off across visual feature depths
Table 3 analyzes hallucination behavior under different visual abstraction depths on POPE-Adversarial. To isolate the effect of feature depth, we mask the inter-layer fusion weights to construct static variants that expose only early (layers 0–7), middle (8–15), or late (16–23) ViT features, as well as a uniform fusion over all layers. We also report the original LLaVA baseline, which relies on a single late-layer representation (layer 2-2).
The results reveal a clear depth-dependent trade-off. Early-layer features are conservative, achieving high precision but low recall, indicating missed positives due to insufficient semantic abstraction. Late-layer features achieve high recall but exhibit a strong “Yes” bias, leading to increased hallucination under adversarial co-occurrence prompts. Mid-layer and uniform fusion partially balance these effects but remain suboptimal.
In contrast, TGIF dynamically adapts the exposed visual abstraction level based on the input query, achieving the best overall F1-score. This confirms that no single fixed depth is optimal for hallucination robustness and motivates text-guided inter-layer fusion.
4.4  Fine-grained perception
@@ 4.5  General reasoning
4.6  Router Layer Selection Dynamics
@@
4.7  Ablation Studies
We conduct ablation studies to analyze the design choices of TGIF, focusing on (i) the effect of multimodal guidance in the router, and (ii) the role of the entropy-based load-balancing loss.
4.7.1  Text-only vs. Multimodal Routing
We compare two routing strategies: text-only router that conditions solely on the question embedding, and a multimodal router that additionally incorporates a global image representation. As shown in Table 1, both variants consistently outperform the LLaVA-1.5 baseline, confirming the effectiveness of dynamic inter-layer fusion.
The multimodal router achieves a substantial gain on HallusionBench(57.31 vs. 49.95), suggesting that global visual context can help the router better distinguish queries requiring semantic reasoning from those demanding conservative verification. It also improves performance on ScienceQA, indicating stronger generalization on complex, multi-step reasoning tasks.
In contrast, the text-only router performs slightly better on POPE and OCRBench, where hallucination often arises from over-reliance on semantic co-occurrence rather than insufficient visual context. In these cases, conditioning routing solely on the linguistic intent of the query encourages more conservative layer selection, reducing sensitivity to visually plausible but absent objects. This behavior aligns with our hypothesis that hallucination is strongly influenced by how visual abstraction depth is exposed to the language model.
Overall, while multimodal routing can benefit tasks that require global semantic understanding, we adopt the text-only router as our final configuration due to its stronger and more stable performance on hallucination-sensitive and fine-grained perception benchmarks. This choice also better isolates the effect of query-driven control over visual representation depth, which is the central focus of TGIF.
4.7.2  Effect of Load-Balancing Coefficient $\lambda$
@@
Figure 5: Qualitative comparison on POPE adversarial. Left: original image. Middle: baseline LLaVA text-conditioned relevance map. Right: TGIF text-conditioned relevance map. The baseline model attends strongly to a semantically misleading object (a fire hydrant resembling a traffic light), resulting in a hallucinated positive answer. In contrast, TGIF suppresses attention to this spurious region and redistributes relevance toward broader visual evidence, leading to a correct negative response.
@@
4.8  Computation and Memory Overhead
We analyze the parameter, latency, and memory overhead introduced by TGIF to validate its lightweight design. The minimal runtime overhead is expected for two reasons. First, the router is a shallow MLP evaluated once per input and contributes negligible computation compared to the frozen vision encoder and the large language model. Second, LLaVA already computes and stores intermediate visual features across layers during the forward pass, so TGIF does not introduce additional vision encoder computation or memory allocation. Instead, it reweights existing representations prior to projection.
Overall, these results confirm that TGIF provides dynamic, query-conditioned control over visual abstraction depth while remaining parameter-efficient and inference-efficient, making it suitable for large-scale deployment without compromising throughput or memory constraints.
Table 6: Computation and memory overhead. Parameter count, inference latency, and peak GPU memory usage. TGIF introduces negligible overhead relative to the LLaVA-1.5 baseline.
Model Params (M) Latency (s) Peak Mem (GB)
LLaVA 7062.9 0.2203 27.16
+ TGIF 7065.0 0.2224 27.16
Overhead +0.03% +0.93% +0.00
4.9  Qualitative comparison
Figure 5 illustrates how text-guided inter-layer fusion alters the visual evidence exposed to the language model. The adversarial query asks whether a traffic light is present. Although the image contains no traffic light, the baseline LLaVA attends strongly to a centrally located fire hydrant whose color and shape are highly correlated with traffic lights in the training distribution. This over-reliance on late, semantically aligned visual features leads the model to hallucinate a positive answer.
In contrast, TGIF dynamically reweights visual representations across layers conditioned on the text query. The resulting relevance map deemphasizes the misleading object and distributes attention more broadly across the scene, suppressing language-driven semantic shortcuts. As a result, TGIF correctly answers “No.” While this is a single example, it qualitatively demonstrates how text-guided layer fusion modulates the balance between visual abstraction and language priors, consistent with our quantitative improvements on hallucination benchmarks.
@@ 5  Conclusion
We presented TGIF, a lightweight text-guided inter-layer fusion framework that dynamically controls which visual representations are exposed to the language model. Our analysis shows that hallucination in multimodal LLMs is closely tied to the depth of visual features used for reasoning, with shallow and deep layers exhibiting complementary failure modes. By adapting the visual abstraction level to the input query, TGIF improves visual grounding and reduces hallucination without modifying the vision encoder, increasing token count, or introducing significant computational overhead. Extensive experiments across hallucination, OCR, and general VQA benchmarks demonstrate that TGIF consistently outperforms strong baselines while preserving overall reasoning capability. These results suggest that query-adaptive control of visual feature depth is a promising direction for building more reliable and trustworthy multimodal language models.
Appendix A Architectural Comparison with Prior Layer Fusion Designs
Figure 6 provides a schematic comparison between commonly used multimodal connectors and the proposed text-guided inter-layer fusion (TGIF).
Figure 6: Comparison of layer fusion designs in MLLMs. (a) MLP Connector: Projects visual tokens from a single, fixed (typically penultimate) vision encoder layer. (b) Dense Connector: Aggregates features from multiple layers via concatenation or downsampling with a static fusion pattern. (c) MMFuser: Retrieves shallow-layer features using deep-layer queries through cross-attention to recover local details. (d) TGIF (ours): Introduces a text-guided router that dynamically reweights visual features across layers based on the input query, enabling adaptive control over visual abstraction depth.
Single-layer projection. Most existing MLLMs, including LLaVA-style architectures, employ a simple MLP connector that projects visual tokens from a single, fixed layer of the vision encoder (typically the penultimate layer) into the LLM. While computationally efficient, this design exposes only late-stage semantic representations to the language model, limiting access to low-level and mid-level visual evidence that is often crucial for fine-grained grounding and hallucination verification.
Static multi-layer fusion. Recent works enrich visual representations by aggregating features from multiple encoder layers. DenseConnector concatenates or downsamples features from selected depths, while MMFuser retrieves shallow-layer details using deep-layer queries via cross-attention. Although these approaches improve visual richness, the fusion strategy remains static: the set of layers and their relative contributions are fixed once chosen and do not adapt to the input query.
Text-guided inter-layer fusion (TGIF). In contrast, TGIF treats the layers of a frozen vision encoder as a pool of specialized visual experts and dynamically selects among them conditioned on the input text. A lightweight router predicts a soft distribution over layers for each query, allowing the model to adaptively control which level of visual abstraction is exposed to the LLM. This enables conservative, low-level verification for hallucination-sensitive queries and higher-level semantic aggregation for descriptive or reasoning-oriented prompts.
Overall, this comparison highlights a key distinction between TGIF and prior fusion methods: rather than statically enriching visual representations, TGIF introduces query-dependent control over visual abstraction depth, directly addressing hallucination arising from mismatched visual evidence.
Appendix B Router Variants: Text-only vs. Multimodal
Figure 7 illustrates the difference between the two routing variants used in TGIF. The text-only router predicts layer weights solely from the question embedding, while the multimodal router additionally incorporates a global image representation (the penultimate-layer [CLS] token), enabling routing to depend on both query intent and coarse image content.
Figure 7: Text-only vs. multimodal router in TGIF. Left: The text-only router takes the question embedding $\mathbf{f}_{\text{text}}$ and outputs layer weights $w_{\text{layer}}$ via an MLP and softmax. Right: The multimodal router concatenates a projected text embedding $\mathbf{f}_{\text{text}}\mathbf{W}_{t}$ and a projected global image embedding $\mathbf{f}_{\text{image}}\mathbf{W}_{v}$ (penultimate-layer [CLS]) to form $\mathbf{f}_{\text{multi}}\in\mathbb{R}^{2D_{p}}$ before predicting $w_{\text{layer}}$.
Appendix C Implementation Details
@@
Appendix D Benchmark Descriptions
D.1  Hallucination Evaluation
POPE [20].
@@ D.2  OCR Evaluation
TextVQA [34].
TextVQA is a benchmark designed to evaluate a model’s ability to read and reason about text embedded in natural images. It contains over 45K questions paired with images rich in scene text (e.g., signs, posters, menus), where answering explicitly requires OCR and text-based reasoning. The dataset features a highly diverse, open-ended answer space, making it a challenging testbed for fine-grained visual grounding and text-centric multimodal understanding.
D.3  General Reasoning and Overall Benchmarks.
ScienceQA [30].
ScienceQA is a large-scale multimodal multiple-choice benchmark designed to evaluate multi-hop scientific reasoning across vision and language. It contains over 21K questions drawn from elementary and high school science curricula, spanning natural, social, and language sciences. Nearly half of the questions include visual context, and most are annotated with grounded lectures and detailed explanations, enabling analysis of both answer accuracy and reasoning quality. ScienceQA is widely used to assess general multimodal reasoning and instruction-following capabilities beyond surface-level perception.
GQA [15].
GQA is a large-scale visual question answering benchmark designed to evaluate compositional and multi-step visual reasoning. It contains over 22M questions grounded in real-world images, each paired with a scene graph encoding objects, attributes, and relations. Questions are generated to require explicit reasoning over these structured representations, reducing language priors and dataset bias. In addition to accuracy, GQA provides fine-grained metrics to assess consistency and plausibility, making it a standard benchmark for general visual reasoning and scene understanding.
MMBench [28]
MMBench is a large-scale, fine-grained benchmark designed to systematically evaluate the perceptual and reasoning abilities of vision–language models. It contains 2,974 multiple-choice questions spanning 20 ability dimensions, organized into a three-level taxonomy covering coarse perception, fine-grained perception, and multimodal reasoning. Instead of task-specific evaluation, MMBench assesses models across diverse capability dimensions, providing more diagnostic feedback. To ensure robust and reproducible evaluation, model outputs are matched to answer choices using an LLM-based evaluator, enabling reliable comparison across open-ended generation models.
Appendix E Additional Quantitative Results
E.1  POPE Breakdown
@@
Table 7: POPE results by subset. Comparison of LLaVA-1.5-7B and LLaVA-1.5-7B+TGIF across different hallucination settings. Best results within each subset are shown in bold.
Subset Method F1 Acc Prec Rec Yes (%)
@@
E.2  HallusionBench Leaderboard
Notably, despite utilizing a significantly smaller language backbone (7B parameters), TGIF (49.94%) outperforms several larger open-source baselines, including the 13B-parameter LLaVA-1.5 (46.94%) and the 12.1B-parameter BLIP2-T5 (48.09%) [25, 18]. It secures the third rank overall, trailing only the closed-source models GPT-4V and Claude 3 [12, 37]. This result shows the efficiency of our text-guided fusion strategy: by dynamically routing to the most relevant visual features, TGIF extracts rich grounding signals from the frozen encoder, effectively mitigating hallucination even against models with nearly double the parameter count.
Table 8: HallusionBench correctness leaderboard. We report Question-pair Accuracy (qAcc), Figure Accuracy (fAcc), and All Accuracy (aAcc). Top-3 models under GPT-4–assisted evaluation are highlighted in bold.
Method Params Eval qAcc $\uparrow$ fAcc $\uparrow$ aAcc $\uparrow$
GPT-4V [12] Human 31.42 44.22 67.58
@@ E.2  HallusionBench Leaderboard
Claude 3 [37] GPT-4 21.76 28.61 56.86
Gemini Pro Vision [38] GPT-4 7.69 8.67 36.85
BLIP2-T5 [18] 12.1B GPT-4 15.16 20.52 48.09
Qwen-VL [3] 9.6B GPT-4 5.93 6.65 39.15
Open-Flamingo [2] 9B GPT-4 6.37 11.27 38.44
MiniGPT-5 [48] 8.2B GPT-4 10.55 9.83 40.30
MiniGPT-4 [50] 8.2B GPT-4 8.79 10.12 35.78
InstructBLIP [8] 8.2B GPT-4 9.45 10.11 45.26
BLIP-2 [18] 8.2B GPT-4 5.05 12.43 40.48
mPLUG-Owl v2 [45] 8.2B GPT-4 13.85 19.94 47.30
mPLUG-Owl v1 [43] 7.2B GPT-4 9.45 10.40 43.93
LRV-Instruction [23] 7.2B GPT-4 8.79 13.01 42.78
LLaVA-1.5 + TGIF (Ours) 7B GPT-4 17.36 23.70 49.94
GIT [40] 0.8B GPT-4 5.27 6.36 34.37
Random Chance GPT-4 15.60 18.21 45.96
E.3  OCRBench Performance
Table 9: Detailed results on OCRBench. Breakdown by sub-task: Text Recognition (Recog.), Scene Text VQA ($VQA^{S}$), Document VQA ($VQA^{D}$), Key Information Extraction (KIE), and Handwritten Math Expression Recognition (HMER). Best results are shown in bold.
@@ E.3  OCRBench Performance
Gemini Pro [38] 215 174 128 134 8 659
GPT-4V [12] 167 163 146 160 9 645
mPLUG-Owl2 [45] 153 153 41 19 0 366
LLaVAR [47] 186 122 25 13 0 346
LLaVA-1.5-13B [25] 176 129 19 7 0 331
LLaVA-1.5-7B [25] 160 117 15 5 0 297
+TGIF (Ours) 162 121 24 6 0 313
mPLUG-Owl [44] 172 104 18 3 0 297
InstructBLIP [9] 168 93 14 1 0 276
BLIP-2 [18] 154 71 10 0 0 235
MiniGPT-4 v2 [49] 124 29 4 0 0 157